OpenStudy (ryanchick):

can i have help on this problem 2(y+2) for y=3

OpenStudy (anonymous):

i think it is 3 +2 * 2

OpenStudy (anonymous):

substitute 3 for y and calculate the final answer

OpenStudy (ryanchick):

okay thanks

OpenStudy (anonymous):

2(3+2) = 2*5 = 10

OpenStudy (anonymous):

yup

OpenStudy (ryanchick):

thank you that helped

OpenStudy (anonymous):

welcome